### v1

@@ 5 Acknowledgments
Figure S.9: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal avarage is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1 model.
@@
Figure S.10: Trend per century in Ekman transport derived from meridional wind stress calculated over period 1995-2100 from historical simulation and SSP5 – 8.5 runs for each upwelling cell in the (a) Canary and (b) Benguela Upwelling region over annual means for the medium and low resolution models. The error bar shows the 90% confidence delimitation.

### v2

@@ 5 Acknowledgments
Fig. S.9: Cross-shore transects of $VT_{\mbox{up-seas}}$ (color-filled black contours) for the Canary (top) and Benguela (bottom) upwelling system, within each sub-region (column). $VT_{\mbox{up-seas}}$ corresponds to an along-shore seasonal average over the entire future period ($2081-2100$). The three-month seasonal average is adapted to each sub-region to only consider months most favourable to upwelling. Contours of potential temperature (light grey) and mixed layer depth (solid white) are superimposed. Line plots above each transect show cross-shore profiles of Sea Surface Temperature (solid black), offshore Ekman transport (solid red)and $VT_{m}$ (blue). The model shown here is the CNRM-CM6-1 model.
@@
Fig. S.10: Trend per century in Ekman transport derived from meridional wind stress calculated over period 1995-2100 from historical simulation and SSP5 – 8.5 runs for each upwelling cell in the (a) Canary and (b) Benguela Upwelling region over annual means for the medium and low resolution models. The error bar shows the 90% confidence delimitation.